# Hierarchical Latent Class Models for Mortality Surveillance Using Partially Verified Verbal Autopsies

### v1

@@ 5 Discussion
In this paper, we introduce a hierarchical latent class model framework for modeling partially verified VA data and estimating the sub-population prevalence of a target cause of death. We describe conditions under which valid inference of the prevalence can be made and propose the novel use of structured priors to borrow information across sub-populations. We demonstrate that our model is able to avoid the bias induced by informative sampling of reference deaths and the structured priors can improve the robustness and interpretability of estimated prevalence.
@@
6 Acknowledgments
We would like to thank Dr. Fatima Marinho and Dr. Luiz Fernando Ferraz for sharing the Brazilian COVID-19 surveillance data and helpful discussion of the results.

### v3

@@ 6 Discussion
In this paper, we introduce a hierarchical latent class model framework for modeling partially verified VA data and estimating the sub-population CSMFs of a target cause of death. We describe conditions under which valid inference of the CSMF can be made and propose the novel use of structured priors to borrow information across sub-populations. We demonstrate that our model is able to avoid the bias induced by informative sampling of reference deaths, and the structured priors can improve the robustness and interpretability of the estimated CSMF.
Our work also highlights the need to carefully consider how training data are collected when deploying predictive models in high-stakes health surveillance tasks. In practice, as long as deaths are selected for verification according to a well-defined study design and protocol, the missing at random assumption can be made plausible by including all relevant factors that influence the verification mechanism in the model. Therefore, careful documentation and consideration of the study design are essential. Failures to account for the data collection process can lead to severely biased results, even with sophisticated models, an aspect of particular relevance as the use of machine learning and artificial intelligence becomes increasingly popular in population health research.
@@
Data Availability Statement
All of the computer codes for this paper are available at https://github.com/YuZoeyZhu/COVID_VA. The repository also includes a reproducible report demonstrating model fitting using synthetic data. The full Brazil COVID-19 dataset can be shared upon request with the permission of the Brazilian Ministry of Health.
Acknowledgments
We would like to thank Dr. Fatima Marinho and Dr. Luiz Fernando Ferraz da Silva for sharing the Brazilian COVID-19 surveillance data and helpful discussion of the results.